pls. help me with this!!!!
A plot of glass is 60 m long and 25 m wide. A sidewalk, 3m wide, surrounds the plot. Find the area of the side walk.
What could the formula be to find the area of the sidewalk?? How could I use the measurement of the plot to find the area of the sidewalk. Pls. help I need an explanation about this. I don't get it..
: pls. help me with this!!!!
A plot of glass is 60 m long and 25 m wide. A sidewalk, 3m wide, surrounds the plot. Find the area of the side walk.
What could the formula be to find the area of the sidewalk?? How could I use the measurement of the plot to find the area of the sidewalk. Pls. help I need an explanation about this. I don't get it..

Answer by nerdybill(1129)

A plot of glass is 60 m long and 25 m wide. A sidewalk, 3m wide, surrounds the plot. Find the area of the side walk.
.
The important thing to do is draw a diagram of the problem. Label all the information provided by the problem.
.
Once you do that, you'll see that you have TWO rectangles. The "outer rectangle" is formed by the outside edge of the sidewalk and the "inner rectangle" is formed by the edge of the grass area.
.
The idea is to:
1. calculate the area of the "inner rectangle"
length = 60 ft
width = 25 ft
area = length*width = 60*25 = 1500 sq ft
.
2. calculate the area of the "outer rectangle"
length = 3+60+3 = 66 ft
width = 3+25+3 = 31 ft
area = length*width = 66*31 = 2046 sq ft
.
3. area of sidewalk = "area of outer rectangle" - "area of inner rectangle"
area of sidewalk = 2046 sq ft - 1500 sq ft = 546 sq ft